APPLICATIONS OF A SINGLE-PULSE BROADBAND FTMW SPECTROMETER: THE DYANMIC ROTA-TIONAL SPECTRUM OF 4-FLUOROBUT-1-YNE.

\underline{BRIAN C. DIAN}, KEVIN O. DOUGLASS, GORDON G. BROWN, SCOTT M. GEYER, and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904..

We perform dynamic rotational spectroscopy on vibrationally excited 4-fluorobut-1-yne using our state-of-the-art, chirpedpulse broadband spectrometer. This new spectrometer allowed us to record 11 GHz of spectrum (7.5-18 GHz) with a single microwave pulse to study the energy-dependence of the isomerization kinetics, dramatically reducing data acquisition times. 4-Fluorobut-1-yne is an asymmetric top found in one of two conformers in a supersonic expansion, with fluorine either trans or gauche with respect to the ethynyl group. The trans- gauche isomerization barrier is approximately 1377 cm^{−1}, and the trans form is more stable by 492 cm^{−1}. Infrared laser excitation was used to prepare 4-fluorobut-1-yne in a region of the potential surface where isomerization can occur ( 3000 cm^{−1}). Using our broadband spectrometer, the rotational spectrum of the infrared excited molecule was recorded from 7.5-18 GHz, covering the ΔJ = 1→2 and ΔJ = 2→3 regions of the spectrum. The overall lineshape of the rotational spectrum was fit with a three-state Bloch model, modified for chemical exchange, yielding an isomerization rate of k_{i}so = 2.7x10^{10} s^{−1}, much slower than the rate predicted by RRKM theory: k_{i}so = 5.9x10^{11} s^{−1}. The single eigenstate rotational spectra recorded with our newly developed spectrometer agree with previously measured single eigenstate rotational spectra, recorded in the acetylenic CH-stretch at 3330 cm^{−1}. Using a combination of double (infrared-microwave) and triple resonance (infrared-microwave-microwave) techniques, single eigenstate rotational spectra were recorded for several IR bands. Infrared excitation of the various CH-stretching bands allows us to observe the isomerization dynamics over a broad energy range (2900-3330 cm^{−1}). Observed changes on the rotational spectrum will be discussed.